OpenStudy (anonymous):

How would you turn this into a logarithmic function? f(x)=2.59e^(.08x)

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

divide both sides by that constant 2.59, then take the logarithm of both sides (to base e)

OpenStudy (unklerhaukus):

can you do the first step?

OpenStudy (anonymous):

No I can't I don't know how.

OpenStudy (unklerhaukus):

f(x)=2.59e^(.08x) \(\dfrac{f(x)}{2.59}\) = \(\dfrac{2.59}{2.59}\) e^(.08x)

OpenStudy (anonymous):

Which equates to this? f(x)/2.59=e^(.08x)

OpenStudy (unklerhaukus):

yeap

OpenStudy (unklerhaukus):

so now now take the natural log of both sides, ln [ f(x)/2.59 ] = \ln [ e^(.08x) ] =

OpenStudy (anonymous):

f(x)=2.59e^(.08x) y= 2.59e^(.08x) y/2.59=e^(.08x) In(y/2.59)=In e^(.08x) In y – In 2.59= In e^(.08x) In y = In e^(.08x) + 0.95 In y = (.08x) + 0.95 This Is what I had before. Was it correct? It seemed wrong so I said I had no Idea how to do it.

OpenStudy (anonymous):

I replaced f(x) with y. to graph with. (suppose to use w/o technology.)

OpenStudy (unklerhaukus):

f(x) =2.59e^(.08x) you set f(x) = y , this is fine y = 2.59e^(.08x) dividing by the constant y/2.59 =e^(.08x) taking the natural log In(y/2.59) =In e^(.08x) log rules , In y – In 2.59 = (.08x) In y = (.08x) + In 2.59 i'd leave answer as this or maybe In f(x) = (.08x) + In 2.59

OpenStudy (unklerhaukus):

oh, you have to graph it? then i guess approximating the log with a decimal is a good idea too then

OpenStudy (unklerhaukus):

you had it right

OpenStudy (anonymous):

Oh wow, thank you. I guess I surprised myself.